# Quantum Frontiers in High Energy Physics

### v1

@@ I Introduction
On the other hand, quantum information science (QIS) has made significant strides. Techniques for controlling single atoms, electrons, and photons have led to innovations in quantum sensors with high sensitivity to weak forces [4]. Advancements in techniques for manipulating groups of these particles hold great promise for the construction of scalable quantum computing platforms, along with the development of new quantum algorithms that demonstrate potential speedups over classical algorithms [5]. These achievements have not only taken us into uncharted territory but have also transformed our approach to scientific inquiry in HEP. In this review, we aim to provide a concise yet comprehensive overview of the areas in HEP that are advanced or could be advanced by the developments of QIS, as outlined in Fig. 1. We will explore how quantum devices, with their unprecedented precision, could detect subtle effects of potential new physics beyond the Standard Model (BSM). Additionally, we will discuss the impact of quantum algorithms and the emergence of future large-scale quantum computers on our theoretical studies of non-perturbative dynamics by solving complex problems in polynomial time, as well as their potential to help uncover new observations by handling the vast and intricate data from HEP experiments. Furthermore, we highlight the importance of incorporating quantum properties into HEP experiments, as this could immediately test our understanding of quantum mechanics at unprecedented high-energy scales and potentially open new avenues to look for hints of BSM from a quantum mechanical perspective.
This lays out the structure of the current review. We dive into recent developments in testing fundamental principles and searching for new physics with quantum sensing and its increased sensitivity in Sec. (II). Then we discuss recent explorations in quantum computing as new paradigms to study dynamics in the non-perturbative regime, including the developments of general framework for QCD calculations and low-dimensional benchmark studies. In Sec. (III), we provide a brief summary of the growing efforts to assess the applicability of quantum machine learning in analyzing HEP data and to explore its potential advantages in enhancing the efficiency of uncovering new observations. Sec. (V) aims to bring recent results in testing quantum properties at HEP experiments into the spotlight, as well as summarize recent new developments in phenomenological studies. We will then discuss perspectives for future directions to explore and conclude in Sec. (VI).
@@ II Quantum Sensing
Quantum technologies are employed to search for such wave-like DM, facilitating unprecedented precision in sensing capabilities. Resonant detectors such as microwave cavities or superconducting circuits are especially well-suited for searching for these forms of DM by matching the resonant frequency to their mass. Axions in the presence of a strong background magnetic field and dark photons, can induce effective currents through their couplings to the electromagnetic fields. Since the experiment sensitivity increases with the quality factor of the resonant detector, a superconducting radio-frequency (SRF) cavity with an exceptionally high-quality factor $Q>10^{10}$ has generated a lot of interest in DM searches [17]. In particular, it can be shown that the scan rate of such a haloscope experiment is proportional to the loaded quality factor even if the cavity bandwidth is much narrower than the DM halo line shape [18]. For example, Ref. [19] has conducted a SRF cavity-based experiment looking for dark photon DM at around 1.3 GHz, placing the world’s leading constraint on the dark photon-photon mixing parameter for this mass range. Quantum sensors such as superconducting qubit [20], trapped electrons [21], Rydberg atoms [22] and NV centers [23], also offer powerful tools for the direct detection of wave-like DM.
@@
Quantum metrology provides various techniques to further improve the detection sensitivity and scan rate of DM detection. We may achieve Distributed Quantum Sensing (DQS) [30, 31] and quantum-enhanced parameter estimation [32] using a quantum network with entangled sensors. In this way, we can use $M$ entangled quantum sensors to achieve an $O(M^{2})$-enhancement in scan rate with a proper set of quantum operations [33, 34, 35, 36, 37, 38]. As an example, Figure 2 shows a DQS setup for cavity-based DM searches, with the dependence of its scan rate on the number of cavities depicted in Figure 3. A non-trivial initial state in the detector with a high occupation number can also increase the signal rate by stimulated emission [39].
@@
Since the first detection in 2015 [123], the ground-based observatories LIGO, Virgo, and KAGRA have observed over 90 gravitational wave events [124], predominantly from binary black hole mergers, but also including binary neutron star and neutron star-black hole systems [125, 126, 127]. Quantum sensing has significantly improved the sensitivity of current detectors. Advanced LIGO and other detectors now routinely use squeezed quantum states of light [128] to reduce quantum noise, particularly shot noise at high frequencies. Additionally, the large mirrors in LIGO, weighing 40 kg, have been shown to be entangled by the squeezed states of light in the interferometer, demonstrating the reach of quantum effects at truly macroscopic scales [129]. The field is also expanding beyond ground-based interferometers. Space-based detectors like the Laser Interferometer Space Antenna (LISA) [130] are under development, aiming to access lower frequency gravitational waves. Meanwhile, atom interferometry [131, 132, 133, 134, 135, 136] is being explored to bridge the mid-frequency gap between ground and space-based detectors.
@@ III Quantum Simulation
The dynamical properties of matter in HEP encompass diverse topics, including non-equilibrium dynamics in the early universe, non-perturbative real-time dynamics that involve parton distribution functions (PDFs) and fragmentation functions (FFs), the phases of matter in heavy ion collisions and in neutron stars, and coherent processes in neutrino physics. All of these stem from the nature of quantum fields upon which our theoretical framework is built. First-principle calculations of these dynamics with classical computers require resources that are exponential functions of the system size, due to the notorious sign problem. Recent experimental advancements in controlling quantum many-body systems, such as those described in [143], have ushered in a new era of computational paradigms: quantum computing. By efficiently exploring vast Hilbert spaces and simulating local Hamiltonians [144], quantum computing offers advantages for performing first-principles calculations of these dynamics in polynomial time. Advances in this direction require lines of research from theory to experiment, including developments of general quantum algorithms to simulate dynamics of QFT on quantum computers, as well as benchmark studies in the noisy intermediate-scale quantum (NISQ) era that could provide guidelines for our understandings of the nature and also hardware developments [145, 146, 147, 148].
@@
After Feynman posed the question of simulating QFT on a quantum computer [149], quantum algorithms have been progressively developed, starting with simulations of the dynamics of quantum systems [144] with a fixed number of particles, extending to lattice gauge theories (LGTs) [150], and later, by Jordan, Lee, and Preskill (JLP) [151], addressing the convergence to the continuum limit. Quantum algorithms developed so far often involve two essential truncations: discretizing the space and digitizing the field, which allow for mapping the field’s infinite degrees of freedom (DoF) onto a finite number of qubits. They further require the preparation of an initial state from the vacuum, the real-time evolution of the system, and the measurement of observables. Theoretical studies are also emerging that investigate the systematic errors from these truncations and approximations used in the real-time evolution, as well as develop efficient methods for achieving the continuum limit.
@@
Discretization – The first essential truncation is to discretize the space on lattices. Other discretization methods on lattices that differ from spatial lattices can be found in, e.g. [152, 153, 154, 155]. The gauge-invariant Hamiltonian to describe the theory on a lattice was developed by John Kogut and Leonard Susskind [156, 157] in the temporal gauge $A_{0}=0$, in which fermions are placed on lattice sites and gauge bosons on the links between sites. This is the well-known KS Hamiltonian, with errors to the Hamiltonian in the continuous space of $\mathcal{O}(a)$, where $a$ is the lattice spacing. The KS Hamiltonian is most commonly considered in quantum simulations at the present time. Including more gauge-invariant terms in the Hamiltonian can improve the convergence rate to the continuous-space theory by pushing the discretization errors to higher orders of $a$ [158, 159, 160, 161, 162, 163]. Quantum simulations of these improved Hamiltonians have recently been investigated [164, 165, 166] with the primary goal of reducing the resource consumption required in approaching the continuum limit. Nonetheless, more effort is needed to quantify improvements in the convergence rate using non-perturbative methods.
@@
Digitization – The second essential truncation involves digitizing the continuous field DoF into finite ones to allow their mapping onto qubits. Since the temporal gauge in the Hamiltonian formulation doesn’t completely fix the gauge, the full Hilbert space spanned by all links on the lattice contains redundant degrees of freedom connected via gauge transformations. These redundancies can be eliminated by projecting onto the gauge-invariant subspace through the imposition of Gauss’s law constraints. These two different Hilbert spaces naturally lead to two classes of digitization, with various methods being proposed for each. One class digitizes the full Hilbert space that includes discrete subgroups in the magnetic basis [167, 168, 169], a finite number of irreducible representations or Schwinger bosons in the electric basis [170, 171, 172, 173, 174, 175, 176], fuzzy gauge theory [177, 178], quantum link models [179, 180, 181, 182, 183]. To perform real-time evolutions, quantum algorithms need to be implemented to prepare gauge-invariant initial states out of the redundant Hilbert space. Such algorithms have recently been realized for digitization methods in the magnetic basis [184], while await to be explored for other digitizations. The second class digitizes only the gauge-invariant states, which includes digitizing the plaquettes [185, 186, 187, 188, 189, 190, 191, 192, 193], states outside the maximal tree in the maximal gauge [194], gauge-invariant bosonic and fermionic operators in the loop-string-hadron (LSH) formulations [195, 196, 197], Fock states in the light-front Hamiltonian [198, 199, 200], local multiplet states [201, 202], and others[203, 204, 205].
@@
Digitizing gauge-invariant states offers advantages in reducing the qubit requirement and avoiding the step of preparing gauge-invariant initial states. However, this process can complicate the Hamiltonian by introducing non-local interactions upon implementing Gauss’s law constraints. This non-locality can naively lead to exponential scaling of the gate count with the volume. To address this issue, operator redefinitions have been employed for compact U(1) gauge theories [191, 206], which effectively reduce the non-locality and break the exponential scaling. Nevertheless, the persistence of exponential scaling remains a challenge for non-Abelian gauge theories. On the other hand, when digitizing the full Hilbert space, the Hamiltonian remains local. The inherent redundancies within the Hilbert space have been leveraged to make strides in quantum error correction and mitigation [207, 208, 209, 210, 211, 212, 213, 214, 215, 216, 217, 218]. With the correctable errors identified, error thresholds can be derived [184] below which keeping gauge redundancies for quantum error corrections can achieve higher fidelity than removing these redundancies, providing important guidance for fault-tolerant quantum simulations of gauge theories.
@@
Continuous limits – To predict quantities that can be used as inputs for experiments, the results from quantum simulations need to be extrapolated to the limit of continuous gauge-boson fields and also to the continuous spacetime limit of infinite volume. In approaching the continuous gauge-boson field limit, a sufficient number of DoF need to be retained in the digitizations, while improved Hamiltonians may be needed for discrete subgroup digitization methods, given that only a finite number of crystal-like subgroups exist for non-Abelian gauge theories [248]. Efforts in extrapolating to the continuous spacetime limit and understanding the systematic uncertainties from finite volume in real-time dynamics remain underdeveloped [167, 249, 250], where progress can be made.
State Preparation – In QFT, the dimension of the Hilbert space grows exponentially with the system size, posing significant challenges for initial states preparation in simulating the non-perturbative aspects of real-time and non-equilibrium dynamics in high-energy physics. In the JLP protocol, the interacting wave packets are created by evolving the spatially well-separated free wave packets with a time-dependent Hamiltonian, where the interactions are turned on adiabatically. However, the adiabatic evolution strains the resources of present-day quantum hardware, limiting the application of this methodology to bound state scatterings. This challenge has motivated alternative proposals, such as directly creating wave packets in the interacting theory [251], employing Haag-Ruelle formalism [252] and using sequency truncations [253] within adiabatic framework, using optical theorem [254], Lehmann-Symanzik-Zimmermann reduction formula [255], stochastic sampling[256, 257, 258] and algorithms based on the quantum eigenvalue transformation for unitary matrices [259] beyond the adiabatic framework.
@@
In the NISQ era, benchmarks can be explored on quantum computers for algorithm development to perform measurements, as well as for understanding physics related to real-time scattering, parton distribution and fragmentation at colliders, entanglement entropy in different phases of matter at finite density, etc. Hadron dynamics are simulated using 112 qubits with the initial hadron wave-packets prepared using the SC-ADAPT-VQE algorithm [278]. Phase shifts of scattering processes estimated from asymptotic state long after the collision processes, are shown to be extractable using information at an early stage of collisions [279], avoiding decoherence and deep circuits in the NISQ era. Quantum algorithms utilizing ancillary qubits to reconstruct $n$-time correlation functions have been proposed [280] and have been applied to linear response theory, PDF calculation in the NJL model [264] on quantum platforms. Parton shower including quantum interference effects can be formulated on quantum computer based on toy models [281, 282] and from first principles [283]. Accessing full final-state wavefunction is in general prohibited, new techniques are being developed to enable practical implementations on near-term quantum computers. Classical shadows [284, 285], which construct an approximate classical description of a quantum state from randomized measurements, allow sampling of the density matrix to reduce computational resources substantially for a wide range of observables, e.g. two-point correlation functions of the ground states[284].

### v2

@@ I Introduction
On the other hand, quantum information science (QIS) has made significant strides. Techniques for controlling single atoms, electrons, and photons have led to innovations in quantum sensors with high sensitivity to weak forces [4, 5, 6]. Advancements in techniques for manipulating groups of these particles hold great promise for the construction of scalable quantum computing platforms, along with the development of new quantum algorithms that demonstrate potential speedups over classical algorithms [7]. These achievements have not only taken us into uncharted territory but have also transformed our approach to scientific inquiry in HEP. In this review, we aim to provide a concise yet comprehensive overview of the areas in HEP that are advanced or could be advanced by the developments of QIS, as outlined in Fig. 1. We will explore how quantum devices, with their unprecedented precision, could detect subtle effects of potential new physics beyond the Standard Model (BSM). Additionally, we will discuss the impact of quantum algorithms and the emergence of future large-scale quantum computers on our theoretical studies of non-perturbative dynamics by solving complex problems in polynomial time, as well as their potential to help uncover new observations by handling the vast and intricate data from HEP experiments. Furthermore, we highlight the importance of incorporating quantum properties into HEP experiments, as this could immediately test our understanding of quantum mechanics at unprecedented high-energy scales and potentially open new avenues to look for hints of BSM from a quantum mechanical perspective.
This lays out the structure of the current review. We dive into recent developments in testing fundamental principles and searching for new physics with quantum sensing and its increased sensitivity in Sec. (II). Then in Sec. (III), we discuss recent explorations in quantum computing as new paradigms to study dynamics in the non-perturbative regime, including the developments of general framework for QCD calculations and low-dimensional benchmark studies. In Sec. (IV), we provide a brief summary of the growing efforts to assess the applicability of quantum machine learning in analyzing HEP data and to explore its potential advantages in enhancing the efficiency of uncovering new observations. Sec. (V) aims to bring recent results in testing quantum properties at HEP experiments into the spotlight, as well as summarize recent new developments in phenomenological studies. We will then discuss perspectives for future directions to explore and conclude in Sec. (VI).
@@ II Quantum Sensing
Quantum technologies are employed to search for such wave-like DM, facilitating unprecedented precision in sensing capabilities. Resonant detectors such as microwave cavities or superconducting circuits are especially well-suited for searching for these forms of DM by matching the resonant frequency to their mass. Axions in the presence of a strong background magnetic field and dark photons, can induce effective currents through their couplings to the electromagnetic fields. Since the experiment sensitivity increases with the quality factor of the resonant detector, a superconducting radio-frequency (SRF) cavity with an exceptionally high-quality factor $Q>10^{10}$ has generated a lot of interest in DM searches [19]. In particular, it can be shown that the scan rate of such a haloscope experiment is proportional to the loaded quality factor even if the cavity bandwidth is much narrower than the DM halo line shape [20]. For example, Ref. [21] has conducted a tunable SRF cavity-based experiment looking for dark photon DM at around 1.3 GHz, placing the world’s leading constraint on the dark photon-photon mixing parameter for this mass range. Quantum sensors such as superconducting qubit [22], trapped electrons [23, 24], Rydberg atoms [25] and NV centers [26], also offer powerful tools for the direct detection of wave-like DM.
@@
Quantum metrology provides various techniques to further improve the detection sensitivity and scan rate of DM detection. We may achieve Distributed Quantum Sensing (DQS) [33, 34] and quantum-enhanced parameter estimation [35] using a quantum network with entangled sensors. In this way, we can use $M$ entangled quantum sensors to achieve an $O(M^{2})$-enhancement in scan rate with a proper set of quantum operations [36, 37, 38, 39, 40, 41, 42]. Figure 2 shows a DQS setup for cavity-based DM searches, with the dependence of its scan rate on the number of cavities depicted in Figure 3. A non-trivial initial state in the detector with a high occupation number can also increase the signal rate by stimulated emission [43].
@@
Since the first detection in 2015 [127], the ground-based observatories LIGO, Virgo, and KAGRA have observed over 90 gravitational wave events [128], predominantly from binary black hole mergers, but also including binary neutron star and neutron star-black hole systems [129, 130, 131]. Quantum sensing has significantly improved the sensitivity of current detectors. Advanced LIGO and other detectors now routinely use squeezed quantum states of light [132] to reduce quantum noise, particularly shot noise at high frequencies. Additionally, the large mirrors in LIGO, weighing 40 kg, have been shown to be entangled by the squeezed states of light in the interferometer, demonstrating the reach of quantum effects at truly macroscopic scales [133]. The field is also expanding beyond ground-based interferometers. Space-based detectors such as LISA [134], Taiji [135] and TianQin [136] are under development, aiming to access lower frequency gravitational waves. Meanwhile, atom interferometry [137, 138, 139, 140, 141, 142] is being explored to bridge the mid-frequency gap between ground and space-based detectors.
@@ III Quantum Simulation
The dynamical properties of matter in HEP encompass diverse topics, including non-equilibrium dynamics in the early universe, non-perturbative real-time dynamics that involve parton distribution functions (PDFs) and fragmentation functions (FFs), the phases of matter in heavy ion collisions and in neutron stars, and coherent processes in neutrino physics. All of these stem from the nature of quantum fields upon which our theoretical framework is built. First-principle calculations of these dynamics with classical computers require resources that are exponential functions of the system size, due to the notorious sign problem. Recent experimental advancements in controlling quantum many-body systems, such as those described in [149], have ushered in a new era of computational paradigms: quantum computing. Though quantum devices are known to be susceptible to errors due to the delicate nature of quantum states, both hardware technologies and theoretical studies in quantum error correction are advancing (see, e.g., [150, 151]) to push the error rates of quantum devices below the thresholds [152], where reliable large-scale (fault-tolerant) quantum simulations can be achieved. By efficiently exploring vast Hilbert spaces and simulating local Hamiltonians [153], quantum computing offers advantages for performing first-principles calculations of these dynamics in polynomial time. Advances in this direction require lines of research from theory to experiment [154, 155, 156, 157, 158], including developments of general quantum algorithms to simulate dynamics of QFT non-perturbatively, as well as benchmark studies in the noisy intermediate-scale quantum (NISQ) era that could provide guidelines for our understandings of the nature and also hardware developments. For explorations of quantum algorithms in perturbative calculations, see e.g. [159, 160, 161].
@@
After Feynman posed the question of simulating QFT on a quantum computer [162], quantum algorithms for non-perturbative calculations have been under development, starting with simulations of the dynamics of quantum systems [153] with a fixed number of particles, extending to lattice gauge theories (LGTs) [163], and later, by Jordan, Lee, and Preskill (JLP) [164], addressing the convergence to the continuum limit. Algorithms developed so far often involve two essential truncations: discretizing the space and digitizing the bosonic field, which allow for mapping the field’s infinite degrees of freedom (DoF) onto a finite number of qubits. They further require preparations of initial states from the vacuum, the real-time evolution of the system, and the measurement of observables. Theoretical studies are emerging that investigate the systematic errors from these truncations and approximations used in the real-time evolution, as well as develop efficient methods for achieving the continuum limit.
@@
Discretization – The first essential truncation is to discretize the space on lattices. Other discretization methods on lattices that differ from spatial lattices can be found in, e.g. [165, 166, 167, 168]. The gauge-invariant Hamiltonian to describe the gauge theories on a lattice was developed by John Kogut and Leonard Susskind [169, 170] in the temporal gauge $A_{0}=0$, in which fermions are placed on lattice sites and gauge bosons on the links between sites. This is the well-known KS Hamiltonian, with errors to the Hamiltonian in the continuous space of $\mathcal{O}(a)$, where $a$ is the lattice spacing. The KS Hamiltonian is most commonly considered in quantum simulations at the present time. Including more gauge-invariant terms in the Hamiltonian can improve the convergence rate to the continuous-space theory by pushing the discretization errors to higher orders of $a$ [171, 172, 173, 174, 175, 176]. Quantum simulations of these improved Hamiltonians have recently been investigated [177, 178, 179] with the primary goal of reducing the resource consumption required in approaching the continuum limit. Nonetheless, more effort is needed to quantify improvements in the convergence rate using non-perturbative methods.
@@
Digitization – The second essential truncation involves digitizing the continuous field DoF into finite ones to allow their mapping onto qubits. Since the temporal gauge in the Hamiltonian formulation doesn’t completely fix the gauge, the full Hilbert space spanned by all links on the lattice contains redundant degrees of freedom connected via gauge transformations. These redundancies can be eliminated by projecting onto the gauge-invariant subspace through the imposition of Gauss’s law constraints. These two different Hilbert spaces naturally lead to two classes of digitization, with various methods being proposed for each. One class digitizes the full Hilbert space that includes discrete subgroups in the magnetic basis [180, 181, 182, 183, 184], a finite number of irreducible representations or Schwinger bosons in the electric basis [185, 186, 187, 188, 189, 190, 191], fuzzy gauge theory [192, 193], quantum link models [194, 195, 196, 197, 198, 199]. To perform real-time evolutions, quantum algorithms need to be implemented to prepare gauge-invariant initial states out of the redundant Hilbert space. Such algorithms have recently been realized for digitization methods in the magnetic basis [200], while await to be explored for other digitizations. The second class digitizes only the gauge-invariant states, which includes digitizing the plaquettes [201, 202, 203, 204, 205, 206, 207, 208, 209], states outside the maximal tree in the maximal gauge [210], gauge-invariant bosonic and fermionic operators in the loop-string-hadron (LSH) formulations [211, 212, 213], Fock states in the light-front Hamiltonian [214, 215, 216, 217, 218, 219], local multiplet states [220, 221], and others[222, 223, 224, 225, 226].
@@
Digitizing gauge-invariant states offers advantages in reducing the qubit requirement and avoiding the step of preparing gauge-invariant initial states. However, this process can complicate the Hamiltonian by introducing non-local interactions upon implementing Gauss’s law constraints. This non-locality can naively lead to exponential scaling of the gate count with the volume. To address this issue, operator redefinitions have been employed for compact U(1) gauge theories [207, 227], which effectively reduce the non-locality and break the exponential scaling. Nevertheless, the persistence of exponential scaling remains a challenge for non-Abelian gauge theories. On the other hand, when digitizing the full Hilbert space, the Hamiltonian remains local. The inherent redundancies within the Hilbert space have been leveraged to make strides in quantum error correction and mitigation [228, 229, 230, 231, 232, 233, 234, 235, 236, 237, 238, 239, 240, 241]. With the correctable errors identified, error thresholds can be derived [200] below which keeping gauge redundancies for quantum error corrections can achieve higher fidelity than removing these redundancies, providing important guidance for fault-tolerant quantum simulations of gauge theories.
@@
Continuous limits – To predict quantities used as inputs for experiments, the results from quantum simulations need to be extrapolated to the limit of continuous gauge-boson fields and also to the continuous spacetime limit of infinite volume. In approaching the continuous gauge-boson field limit, a sufficient number of DoF need to be retained in the digitizations, while improved Hamiltonians may be needed for discrete subgroup digitization methods, given that only a finite number of crystal-like subgroups exist for non-Abelian gauge theories [271]. Efforts in extrapolating to the continuous spacetime limit and understanding the systematic uncertainties from finite volume in real-time dynamics remain underdeveloped [181, 272, 273], where progress can be made.
State Preparation – In QFT, the dimension of the Hilbert space grows exponentially with the system size, posing significant challenges for initial states preparation in simulating non-perturbative aspects of real-time and non-equilibrium dynamics in HEP. In the JLP protocol, the interacting wave packets are created by evolving the spatially well-separated free wave packets with a time-dependent Hamiltonian, where the interactions are turned on adiabatically. However, the adiabatic evolution strains the resources of present-day quantum hardware, limiting the application of this methodology to bound state scatterings. This challenge has motivated alternative proposals, such as directly creating wave packets in the interacting theory [274], employing Haag-Ruelle formalism [275] and using sequency truncations [276] within adiabatic framework, using optical theorem [277], Lehmann-Symanzik-Zimmermann reduction formula [278], stochastic sampling[279, 280, 281] and algorithms based on the quantum eigenvalue transformation for unitary matrices [282] beyond the adiabatic framework.
@@
In the NISQ era, benchmarks can be explored on quantum computers for algorithm development to perform measurements, as well as for understanding physics related to real-time scattering, parton distribution and fragmentation at colliders, entanglement entropy in different phases of matter at finite density, etc. Hadron dynamics are simulated using 112 qubits with the initial hadron wave-packets prepared using the SC-ADAPT-VQE algorithm [301]. Phase shifts of scattering processes estimated from asymptotic state long after the collision processes, are shown to be extractable using information at an early stage of collisions [302], avoiding decoherence and deep circuits in the NISQ era. Quantum algorithms utilizing ancillary qubits to reconstruct $n$-time correlation functions have been proposed [303] and have been applied to linear response theory, PDF calculation in the NJL model [287] on quantum platforms. Parton shower including quantum interference effects can be formulated on quantum computer based on toy models [304, 305] and from first principles [306]. Thermalization dynamics of $U(1)$ gauge theory in 1+1D has been studied on quantum computers using the quantum link models [307]. Accessing full final-state wavefunction is in general prohibited, new techniques are being developed to enable practical implementations on near-term quantum computers. Classical shadows [308, 309], which construct an approximate classical description of a quantum state from randomized measurements, allow sampling of the density matrix to reduce computational resources substantially for a wide range of observables, e.g. two-point correlation functions of the ground states[308].